OpenStudy (anonymous):

Which equation represents a line that is parallel to the line y=-3x+4

OpenStudy (anonymous):

Do you have options?

OpenStudy (anonymous):

y=3x+1 y=-2x+4 y=-3x+1 y=-1/3x+2

OpenStudy (anonymous):

C. Y = -3x + 1